# MaskMoE: Boosting Token-Level Learning via Routing Mask in Mixture-of-Experts

### v2

@@ Abstract
Scaling the size of a model enhances its capabilities but significantly increases computation complexity. Mixture-of-Experts models (MoE) address the issue by allowing model size to scale up without substantially increasing training or inference costs. Despite their promising results, MoE models encounter several challenges. Primarily, for dynamic routing methods, the dispersion of training tokens across multiple experts can lead to underfitting, particularly for infrequent tokens. Additionally, while fixed routing methods can mitigate that issue, they compromise on the diversity of representations. In this paper, we propose MaskMoE, a method designed to enhance token-level learning by employing a routing masking technique within the Mixture-of-Experts model. MaskMoE is capable of maintaining representation diversity while achieving more comprehensive training. Experimental results demonstrate that our method outperforms previous dominant Mixture-of-Experts models in terms of both perplexity (PPL) and downstream task performance.
@@ 1 Introduction
Large language models have achieved promising performance on various downstream natural language tasks Touvron et al. (2023a); Dai et al. (2022); Brown et al. (2020); Anil et al. (2023). Moreover, according to the scaling law  Kaplan et al. (2020); Hoffmann et al. (2022), as the model size increases, the model’s size will continue to grow. However, for dense language models, the computational costs of continuing to scale up are substantial. In order to further scale up models within computational budgets, sparse activation networks Child et al. (2019); Du et al. (2022a) receive widespread attention due to their ability to significantly reduce computational costs by using only a part of parameters per input. A widely studied approach is Mixture-of-Experts (MoE)  Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024); Fedus et al. (2022); Roller et al. (2021), which trains multiple expert networks but selects only a subset to process specific inputs. Compared to dense networks of the same model size, MoE effectively reduces computational costs and achieves better results in both PPL and downstream tasks Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024).
@@
In MoE models, for each input (e.g., a token in the language models), the router in front of the experts needs to decide which tokens to distribute to which experts. The current commonly used dynamic routing method in MoE is to select the expert with the top $k$ highest confidence based on a probability distribution output by an intermediate layer with learnable parameters. Previous works Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024); Fedus et al. (2022) show that MoE models trained based on dynamic routing achieve better performance than dense models with the same amount of training computation.
However, this dynamic routing method introduces a new challenge: the routing fluctuation problem Dai et al. (2022). It implies that as training moves on, the same token is assigned to different experts in different iterations. This has two negative impacts. Firstly, each expert is trained on only a subset of the occurrences of an identical token, resulting in each expert not learning enough about this token. As previous work Shazeer et al. (2017) has indicated, when there are too many experts, the MoE model exhibits underfitting, and even shows higher PPL than its counterparts with fewer experts. The routing fluctuation problem has relatively less impact on the learning of frequent tokens, as those tokens have sufficient occurrences to ensure that each expert receives adequate training. On the contrary, for infrequent tokens, the routing fluctuation problem can disperse each of them across various experts, which can lead to underfitting for them. Secondly, the knowledge each expert learns from the corresponding occurrence subset w.r.t an identical token is hard to share among experts. Such lack of knowledge sharing and interaction among experts may negatively impact the model’s performance Zhao et al. (2024). The above two issues hinder scaling the number of experts for an MoE model. Because, as the number of experts increases while keeping the training data constant, the average number of tokens assigned to each expert further decreases, exacerbating the above two issues.
@@
Figure 1: Illustration of random routing masking method. For each token, some experts will be randomly masked, with only a subset of experts being visible, e.g., for the token “in” only expert FFN #1, #2, and #3 are visible.
@@
To alleviate the issue of routing fluctuations, the previously proposed fixed routing method Roller et al. (2021); Dai et al. (2022) is a solution. Specifically, Hash Layer Roller et al. (2021) pre-assigns the same token to a fixed routing expert. For instance, as shown in Figure 1, the token “deliberate” can only be routed to the expert FFN#2. The fixed routing method to send all the occurrences of a token to the same expert is advantageous for the token’s thorough learning, especially for infrequent tokens. However, fixed routing based on the Hash Layer represents a deficiency in diversity, especially for those frequent tokens. Previous works He et al. (2023); Yang et al. (2019) show that increasing the number of optional experts for all occurrences of an identical can enhance the diversity of representations. Frequent tokens probably require to be encoded by more experts due to their broader range of usage Koranda et al. (2018). For instance, the token “the” appears in various contexts and owns different semantic meanings, and having a more diverse representation helps to distinguish it. Therefore, based on the observations above, we need a fixed routing strategy for infrequent tokens, while also needing more experts for frequent ones.
@@
To meet the requirements above. In this paper, we propose a routing method called MaskMoE, which includes a routing masking method that adjusts the number of visible experts for tokens of different frequencies, by generating a masking vector for each token in the vocabulary before training. Specifically, for each infrequent token, the proposed MaskMoE employs routing masking to retain only one visible expert to which the token can be routed, in each MoE layer. For each frequent token, the proposed MaskMoE has a lower routing masking rate to allow more visible experts for it in each MoE layer. MaskMoE can enhance token-level learning through such a routing mask. Specifically, MaskMoE makes the model learn more intensively about infrequent tokens by routing each to an identical expert in each MoE layer, enabling the expert to be fully trained w.r.t that token. Meanwhile, MaskMoE maintains the diversity of representations for frequent tokens by still having multiple experts available for routing, which benefits their representation diversity He et al. (2023); Yang et al. (2019). Despite frequent tokens being routed to multiple experts, their sufficient quantities can still ensure thorough training. Our experimental results indicate that MaskMoE outperforms previous MoE models, in terms of either PPL or downstream task performance.
@@
  • We validate the effectiveness of the proposed MaskMoE with extensive experiments. Experimental results show that it consistently outperforms previous Mixture-of-Experts models.
@@ 2.1 Language Models
Language models are statistical models designed to optimize the likelihood of token sequences in training data Touvron et al. (2023a). Initially, language models relied on $n$-gram statistics Bahl et al. (1983); Katz (1987); Kneser and Ney (1995). Subsequently, the emphasis shifted to neural network-based models, particularly Recurrent Neural Networks Mikolov et al. (2010) and their variants like LSTMs Graves (2013). Those models have demonstrated the ability to learn intricate patterns within textual data, achieving significant success in various language modeling applications.
In recent times, Transformers have become the predominant architecture for language models. Notable examples include BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), GPT-2 Radford et al. (2019), UniLM Dong et al. (2019), and T5 Raffel et al. (2020). The introduction of GPT-3 Brown et al. (2020), which boasts 175 billion parameters, marked a significant milestone due to its exceptional performance across numerous downstream tasks. And that then led to a surge in research focusing on large generative language models, including prominent works such as Gopher Rae et al. (2021), PaLM Chowdhery et al. (2022), Pythia Biderman et al. (2023), OPT Zhang et al. (2022), GLM Du et al. (2022b); Zeng et al. (2023) and LLaMA Touvron et al. (2023a, b). Currently, GPT4 OpenAI (2023), trained with Reinforcement Learning from Human Feedback (RLHF), achieves truly remarkable results.
@@ 3 Method
In this section, we first review the training of the MoE models. Then, we elaborate on the proposed MaskMoE. Finally, we introduce the load balance loss.
@@ 3.1 Reviewing the Mixture-of-Experts
The routing of MoE can be divided into fixed routing Roller et al. (2021); Dai et al. (2022) and dynamic routing with learnable parameters Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024); Fedus et al. (2022). The common method for fixed routing is to use random hashing to determine Roller et al. (2021), before training, which experts each token will be sent to. In contrast, dynamic routing employs a routing layer with learnable parameters to decide which expert should process an input token.
In terms of structure, the FFN layer of a dense language model is generally replaced with sparse expert modules to build an MoE model. Given a tokenized input sequence $\mathbf{x}={x_{1},x_{2},...,x_{T}}$ consisting of $T$ tokens , and then generates a probability distribution $\mathbf{p}_{v}$ over the vocabulary as output. In nearly all implementations, the Cross-Entropy loss is used as the loss function to maximize the predicted probability $\mathbf{p}_{v}^{x_{i}}$ w.r.t the ground-truth token $x_{i}$. The training loss $\mathcal{L}_{lm}$ of the generative language model can be formulated as follows.
$\mathcal{L}_{lm}$ $=-\sum_{i=2}^{T}\log(\mathbf{p}_{v}^{x_{i}})$ (1)
$s.t.,\mathbf{p}_{v}$ $=\text{softmax}(W\mathbf{H}^{L})$ (2)
$\mathbf{H^{L}}$ $=\text{Transformer}(x_{1}^{0},x_{2}^{0},\ldots,x_{T}^{0})$ (3)
Here, $\mathbf{H^{L}}=[\mathbf{h}_{1}^{L},\mathbf{h}_{2}^{L},\ldots,\mathbf{h}_{T}^{L}]$ denotes the hidden states of the last layer. With $\mathbf{H}^{L}$, a linear projection layer $W$ is introduced to derive the predicted probability distribution $\mathbf{p}_{v}$ over the vocabulary, where $\mathbf{p}_{v}=[\mathbf{p}^{x_{1}}_{v},\mathbf{p}^{x_{2}}_{v},\ldots,\mathbf{p}^{x_{T}}_{v}]$. The Transformer model consists of $L$ Transformer layers. Each layer is composed of a self-attention sublayer and an FFN sublayer, where FFN is a two-layer fully connected network.
$\mathbf{\hat{h}}_{1}^{l},\mathbf{\hat{h}}_{2}^{l},\ldots,\mathbf{\hat{h}}_{T}^{l}=SelfAttn(\mathbf{h}_{1}^{l-1},\mathbf{h}_{2}^{l-1},\ldots,\mathbf{h}_{T}^{l-1})$ (4)
$\mathbf{h}_{1}^{l},\mathbf{h}_{2}^{l},\ldots,\mathbf{h}_{T}^{l}=FFN(\mathbf{\hat{h}}_{1}^{l},\mathbf{\hat{h}}_{2}^{l},\ldots,\mathbf{\hat{h}}_{T}^{l})$ (5)
$l$ represents the $l$-th transformer layer. MoE methods replace the FFN sub-layer with the MoE module. A MoE model can replace the FFN with a MoE module for either a single layer, several layers, or even all layers. Each MoE module consists of multiple FFNs, whose outputs are mixed with the gating function $P(\cdot)$, represented by the following formula.
$\mathbf{h}_{t}^{l}=\sum_{i}^{N}P_{i}(\mathbf{\hat{h}}_{t}^{l})\cdot FFN_{i}(\mathbf{\hat{h}}_{t}^{l})$ (6)
$s.t.|P(\mathbf{\hat{h}}_{t}^{l})|_{0}<<N$ (7)
Where $N$ represents the number of experts in a single MoE layer. The majority of elements in the distribution of a gated $P(\cdot)$ output are zeros, i.e., only a small portion of experts will be activated. Therefore, increasing the total number of experts in MoE does not increase computation or inference time.
@@ 3.2 Proposed MaskMoE
Dynamic routing uses a learnable method where tokens are assigned to different experts during training, leading to routing fluctuations. On the other hand, fixed routing, like hash layers, assigns each token to a specific expert, which can also limit representation diversity. To mitigate the routing fluctuations and the diversity of representations in MoE models, we propose MaskMoE to boost token-level learning via routing mask, as illustrated in Figure 1. For each token, a fixed random masking is applied to the experts, allowing them to be routed only to their respective subsets of experts. The formula is represented as follows.
$\mathbf{p}$ $=\text{Softmax}(W_{g}\mathbf{\hat{h}_{t}^{l}}+\mathbf{m})$ (8)
Where $\mathbf{m}$ is the masking vector used to control the visibility of token-specific experts. For visible experts, the corresponding elements in $\mathbf{m}$ are 0, while for invisible experts, the corresponding elements in $\mathbf{m}$ are $\infty$- ∞. The masking vector is determined before training and does not change during the training process. For multi-layer MoE models, we reuse the same masking vector $\mathbf{m}$ across different MoE layers. Here, $W_{g}$ is the parameter of the gating function. The initial energy masking vector for a token $t$ can be represented by the following formula.
@@
$\mathbf{C}$ =$=\{\text{C}_{i}\}_{i=1}^{V}\sim\mathcal{U}(\{0:N-1\})$
@@
We denote the number of visible experts as $V$. $N$ represents the number of experts in a single MoE layer. $\mathcal{U}(\{0:N-1\})$ denotes a uniform distribution over the set of integers from 00 to $N-1$. $\text{C}_{i}$ denotes the index of the selected expert. It is worth noting that for infrequent tokens, each token has only one visible expert, i.e., $V=1$. This promotes more thorough learning by the selected expert for such infrequent tokens. For frequent tokens, the number of visible experts is greater than 1, i.e., $1<V<=N$, and note that it is common for $V$ to be less than $N$. We believe that an appropriately sized $V$ not only allows frequent tokens to have higher representation diversity compared to fixed routing but also enables more thorough training for an identical token compared to dynamic routing. The complete process of MaskMoE is illustrated in Algorithm 1.
Algorithm 1 Proposed MaskMoE Algorithm
0:  Training dataset after tokenization $D$, the vocabulary of the tokenizer $T_{v}$, maximal training step number $S$, the number of visible experts $V$, the number of experts $N$
1:  Initialize the routing mask matrix, $\mathbf{M}\leftarrow-\infty$
2:  for each $t$ in $T_{v}$ do
3:     $\mathbf{C}=\{\text{C}_{i}\}_{i=1}^{V}\sim\mathcal{U}(\{0:N-1\})$
4:     $\mathbf{m}^{t}=\mathbf{M}(t)$ The corresponding column w.r.t $t$
5:     $\mathbf{m}_{j}^{t}=0,\quad\forall j\in\mathbf{C}$
6:  end for
7:  for $i=1$ to $S$ do
8:     Load an input sequence $\mathbf{x}$ from $D$
9:     # Load the routing mask for each token in $\mathbf{x}$
10:     $\mathbf{m}=\mathbf{M}(\mathbf{x})$
11:     Train the MoE models with $\mathbf{x}$ as the input, and add $\mathbf{m}$ to the router according to Equation 10
12:  end for
3.3 Load Balance Loss
MoE models typically require distributed training, where different experts are deployed across various computation nodes (e.g., GPU server). Distributed training can lead to load imbalance, where a minority of expert nodes handle the majority of tokens while others remain idle most of the time, and it can negatively impact training efficiency. It is generally desirable for the number of tokens processed by different experts to be roughly equal. To achieve that, a load balancing loss is commonly introduced in the training of MoE models. We follow previous works Huang et al. (2024); Fedus et al. (2022) and adopt a widely used loss as follows.
$\mathcal{L}_{bal}$ $=N*\sum_{i=1}^{N}w_{i}\cdot R_{i}$ (10)
[…]
$R_{i}$ $=\frac{1}{T}\sum_{j=1}^{T}\mathbf{p}_{i}^{j}$ (12)
Where $T$ represents the number of tokens in a mini-batch. $\mathbf{p}^{j}$ denotes the probability distribution of the routing output for the $j$-th token, while $\mathbf{p}_{i}^{j}$ represents the specific probability value of the $i$-th expert. It is noteworthy that for infrequent tokens, which have only one visible expert, the load balancing loss does not apply to them. Since infrequent tokens are assigned to experts through a uniform random hashing, the overall load is supposed to be natively balanced, thus having no need for the balancing loss. The balancing loss primarily regulates the routing of frequent tokens.
Our final loss is a combination of the language model loss and load-balance loss:
$\mathcal{L}=\mathcal{L}_{lm}+\mathcal{L}_{bal}$ (13)
4 Experiments
Model Configuration Params Activated Params Pile PPL

### v4

@@ Abstract
Scaling the size of a model enhances its capabilities but significantly increases computation complexity. Mixture-of-Experts models (MoE) address the issue by allowing model size to scale up without substantially increasing training or inference costs. In MoE, there is an important module called the router, which is used to distribute each token to the experts. Currently, the mainstream routing methods include dynamic routing and fixed routing. Despite their promising results, MoE models encounter several challenges. Primarily, for dynamic routing methods, the dispersion of training tokens across multiple experts can lead to underfitting, particularly for infrequent tokens. Additionally, though fixed routing methods can mitigate that issue, they compromise on the diversity of representations. In this paper, we propose MaskMoE, a method designed to enhance token-level learning by employing a routing masking technique within the Mixture-of-Experts model. MaskMoE is capable of maintaining representation diversity while achieving more comprehensive training. Experimental results demonstrate that our method outperforms previous dominant Mixture-of-Experts models in terms of both perplexity (PPL) and downstream task performance.
@@ 1 Introduction
Large language models have achieved promising performance on various downstream natural language tasks Touvron et al. (2023a); Dai et al. (2022); Brown et al. (2020); Anil et al. (2023); Chowdhery et al. (2022); Radford et al. (2019); Rae et al. (2021); Biderman et al. (2023). Moreover, according to the scaling law  Kaplan et al. (2020); Hoffmann et al. (2022), as the model size increases, the model’s capabilities will continue to grow. However, for dense language models, the computational costs of continuing to scale up are excessive. In order to further scale up models within computational budgets, sparse activation networks Child et al. (2019); Du et al. (2022a) receive widespread attention due to their ability to significantly reduce computational costs by using only a part of parameters per input. A widely studied approach is Mixture-of-Experts (MoE)  Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024); Fedus et al. (2022); Roller et al. (2021), which trains multiple expert layers but selects only a subset to process specific inputs. Compared to dense networks of the same model size, MoE effectively reduces computational costs and achieves comparable results in both PPL and downstream Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024).
@@
In MoE models, for each input (e.g., a token in the language models), the router in front of the experts needs to decide which experts to feed it in. The commonly used dynamic routing method in MoE is to select the expert with the top $k$ highest confidence based on a probability distribution output by an intermediate layer with learnable parameters that act as the router. Previous works Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024); Fedus et al. (2022) show that MoE models trained based on dynamic routing achieve better performance than dense models with the same amount of training computation.
@@
Figure 1: Illustration of random routing masking method. For each token, some experts will be randomly masked, with only a subset of experts being visible.
However, the dynamic routing method introduces a new challenge: the routing fluctuation problem Dai et al. (2022). It implies that as training moves on, the same token is assigned to different experts in different iterations due to the variation of learnable parameters in the router. That has two negative impacts. a) Each expert is trained on only a subset of the occurrences of an identical token, resulting in each expert possibly not learning enough about the token. As  Shazeer et al. (2017) has indicated, when there are too many experts, the MoE model exhibits underfitting, and even shows higher PPL than its counterparts with fewer experts. The routing fluctuation problem has relatively less impact on the learning of frequent tokens, as those tokens have sufficient occurrences to ensure that each expert receives adequate training. On the contrary, for infrequent tokens, the routing fluctuation problem can disperse each of them across various experts, which probably leads to underfitting for them. b) The knowledge each expert learns from the corresponding occurrence subset w.r.t an identical token is hard to share among experts. Then the underfitting of experts for some tokens, together with the lack of knowledge sharing among experts, can negatively impact the model’s performance Zhao et al. (2024). Both issues above hinder scaling the number of experts for an MoE model. Because, as the number of experts increases while keeping the training data unchanged, the average number of token occurrences assigned to each expert further decreases, exacerbating the issues.
@@
To alleviate the issue of routing fluctuations, fixed routing method Roller et al. (2021); Dai et al. (2022) is proposed, e.g., Hash Layer. Specifically, Hash Layer Roller et al. (2021) pre-assigns each token to a fixed expert. For instance, as shown in Figure 1, all occurrences of the token “deliberate” can only be routed to the expert FFN#2. The fixed routing method to send all the occurrences of a token to the same expert is advantageous for the token’s thorough learning, especially for infrequent tokens. However, fixed routing based on the Hash Layer represents a deficiency in learning representation diversity, especially for those frequent tokens. Previous works He et al. (2023); Yang et al. (2019) show that increasing the number of optional experts for all occurrences of an identical token can help the diversity of learned representations. And frequent tokens probably require to be encoded by more experts due to their broader range of usage Koranda et al. (2018). For instance, the token “play” appears in various contexts and owns different semantic meanings, and thus having a more diverse representation helps to distinguish them.
@@
Therefore, based on the observations above, we need a fixed routing strategy to alleviate the underfitting problem for infrequent tokens, while also requiring more experts for frequent ones to maintain the representational diversity. To meet the requirements above, in this paper, we propose a routing method called MaskMoE, which includes a routing masking method that adjusts the number of visible experts for tokens of different frequencies, by generating a masking vector for each token in the vocabulary before training. Specifically, for each infrequent token, the proposed MaskMoE employs routing masking to retain only one visible expert to which the token can be routed, in each MoE layer. For each frequent token, the proposed MaskMoE allow more visible experts for it in each MoE layer. MaskMoE can enhance token-level learning through such a routing mask. Specifically, MaskMoE makes the model learn more intensively about infrequent tokens by routing each to an identical expert in every MoE layer, enabling the expert to be adequately trained w.r.t that token. Meanwhile, MaskMoE maintains the diversity of representations for frequent tokens by still having multiple experts available for routing, which benefits their representation diversity He et al. (2023); Yang et al. (2019). Despite frequent tokens being routed to multiple experts, their sufficient quantities can still ensure thorough training. Our experimental results indicate that MaskMoE outperforms previous MoE models, in terms of either PPL or downstream task performance.
@@
  • We validate the effectiveness of the proposed MaskMoE with extensive experiments. Experimental results show that it consistently outperforms previous dynamic routing and fixed routing methods for MoE models.
@@ 2.1 Language Models
Language models are statistical models designed to optimize the likelihood of token sequences in training data Touvron et al. (2023a). Initially, language models relied on $n$-gram statistics Bahl et al. (1983); Katz (1987); Kneser and Ney (1995). Subsequently, the emphasis shifted to neural network-based models, particularly Recurrent Neural Networks Mikolov et al. (2010) and their variants like LSTMs Graves (2013). Those models have demonstrated the ability to learn intricate patterns within textual data, achieving significant success in various language modeling applications. In recent times, Transformers have become the predominant architecture for language models. Notable examples include BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), GPT-2 Radford et al. (2019), UniLM Dong et al. (2019), and T5 Raffel et al. (2020). The introduction of GPT-3 Brown et al. (2020), which boasts 175 billion parameters, marked a significant milestone due to its exceptional performance across numerous downstream tasks. And that then led to a surge in research focusing on large generative language models, including prominent works such as Gopher Rae et al. (2021), PaLM Chowdhery et al. (2022), Pythia Biderman et al. (2023), OPT Zhang et al. (2022), GLM Du et al. (2022b); Zeng et al. (2023) and LLaMA Touvron et al. (2023a, b). Currently, GPT4 OpenAI (2023) achieves truly remarkable results.
@@ 3 Method
3.1 Reviewing the Language Models
Given a tokenized input sequence $\mathbf{x}=({x_{1},x_{2},...,x_{T}})$ consisting of $T$ tokens, a language model generates a probability distribution $\mathbf{p}$ over the vocabulary as output for each token. In nearly all implementations, the Cross-Entropy loss is used as the loss function to maximize the predicted probability $\mathbf{P}_{\cdot,i}^{x_{i}}$ w.r.t the ground-truth token $x_{i}$. The training loss $\mathcal{L}_{lm}$ of the generative language model can be formulated as follows:
$\mathcal{L}_{lm}$ $=-\sum_{i=2}^{T}\log(\mathbf{P}_{\cdot,i}^{x_{i}})$ (1)
$s.t.,\quad\mathbf{P}_{\cdot,i}$ $=\text{softmax}(W\mathbf{H}^{L}_{\cdot,i})$
$\mathbf{H}^{L}$ $=\text{Transformer}(x_{1},x_{2},\ldots,x_{T})$
where $\mathbf{P}_{\cdot,i}$ and $\mathbf{H}_{\cdot,i}^{L}$ denote the $i$-th column of the matrix $\mathbf{P}$ and $\mathbf{H}^{L}$. Here, $\mathbf{H}^{L}=[\mathbf{h}_{1}^{L},\mathbf{h}_{2}^{L},\ldots,\mathbf{h}_{T}^{L}]$ denotes the hidden states of the last layer. With $\mathbf{H}^{L}_{\cdot,i}$, a linear projection layer $W$ is introduced to derive the predicted probability distribution $\mathbf{P}_{\cdot,i}$ over the vocabulary. The Transformer model consists of $L$ Transformer blocks. Each block is composed of a multi-head self-attention (MHA) module and an FFN module, where the FFN module is generally a two-layer fully connected network. Formally,
$\mathbf{\hat{h}}_{1}^{l},\mathbf{\hat{h}}_{2}^{l},\ldots,\mathbf{\hat{h}}_{T}^{l}=\text{MHA}(\mathbf{h}_{1}^{l-1},\mathbf{h}_{2}^{l-1},\ldots,\mathbf{h}_{T}^{l-1})$ (2)
$\mathbf{h}_{1}^{l},\mathbf{h}_{2}^{l},\ldots,\mathbf{h}_{T}^{l}=\text{FFN}(\mathbf{\hat{h}}_{1}^{l},\mathbf{\hat{h}}_{2}^{l},\ldots,\mathbf{\hat{h}}_{T}^{l})$ (3)
where $l$ represents the $l$-th Transformer block.
@@ 3.2 Reviewing the Mixture-of-Experts
MoE methods generally replace a single FFN module or multiple/all FFN modules as MoE modules. Each MoE module consists of multiple FFNs, whose outputs are mixed with the routine function $\mathbf{r}(\cdot)$, i.e., the router, as shown in the following formula:
$\mathbf{h}_{t}^{l}=\sum_{i}^{N}\mathbf{r}_{i}(\mathbf{\hat{h}}_{t}^{l})\cdot\text{FFN}_{i}(\mathbf{\hat{h}}_{t}^{l})\qquad s.t.,|\mathbf{r}(\mathbf{\hat{h}}_{t}^{l})|_{0}\ll N$ (4)
where $N$ represents the number of experts in a single MoE module, $\mathbf{r}_{i}$ denotes the routine result w.r.t the $i$-th expert, and $|\cdot|_{0}$ denotes the $L_{0}$-norm. The majority of elements $\mathbf{r}(\cdot)$ are zeros, and thus only a small portion of experts will be activated. Therefore, increasing the total number of experts in MoE does not increase computation or inference time substantially. Regarding the router $\mathbf{r}$ of MoE, it can be divided into fixed routing  Roller et al. (2021); Dai et al. (2022) and dynamic routing with learnable parameters  Fedus et al. (2022); Lepikhin et al. (2021); DeepSeek-AI et al. (2024); Dai et al. (2024). The former commonly use random hashing to determine  Roller et al. (2021), before training, which experts each token will be sent to, i.e., $\mathbf{r}$ is unlearnable and preset. In contrast, the latter employs a routing layer with learnable parameters to decide which expert should process an input token, i.e., $\mathbf{r}$ is to be learned during model training.
@@ 3.3 Proposed MaskMoE
As mentioned before, dynamic routing can cause routing fluctuations, while fixed routing, like hash layers, can limit representation diversity. To mitigate the above two issues in MoE models, we propose MaskMoE to boost token-level learning via routing mask, as illustrated in Figure 1. Specifically, for each token, a fixed random masking is applied to the experts, allowing them to be routed only to their corresponding subsets of experts. The formula is represented as follows:
$\mathbf{r}$ $=\text{softmax}(W_{r}\mathbf{\hat{h}_{t}}^{l}+\mathbf{m}^{t})$ (5)
where $\mathbf{m}^{t}$ is the masking vector used to control the visibility of token-specific experts for the token $t$. For visible experts, the corresponding elements in $\mathbf{m}^{t}$ are 0, while for invisible experts, the corresponding elements in $\mathbf{m}^{t}$ are $\infty$- ∞. The masking vector is determined before training and does not change during the training process. For multi-layer MoE models, we reuse the same masking vector $\mathbf{m}^{t}$ across different MoE layers. Here, $W_{r}$ is the learnable parameter of the routing function. It can be seen that we use the preset $\mathbf{m}^{t}$ to control the visibility of experts like the fixed routing method and meanwhile reserve the learnable parameter $W_{r}$ as the dynamic routing method, and thus we combine the strengths of both. The initial masking vector for a token $t$ can be represented by the following formula:
@@
$\mathbf{C}$ =$=\{\text{C}_{i}\}_{i=1}^{V}\sim\mathcal{U}(\{1:N\})$
@@
where the number of visible experts is $V$. $\mathcal{U}(\{1:N\})$ denotes a uniform distribution over the set of integers from $1$ to $N$, and $\text{C}_{i}$ denotes the index of the selected expert for the $i$-th random sampling. It is worth noting that for infrequent tokens, each token usually has only one visible expert, i.e., $V=1$. That promotes more thorough learning by the selected expert for such infrequent tokens. For frequent tokens, the number of visible experts is generally greater than 1, i.e., $1<V<=N$, and it is common for $V$ to be less than $N$. We believe that an appropriately sized $V$ not only allows frequent tokens to have higher representation diversity compared to fixed routing but also enables more thorough training for an identical token compared to dynamic routing.
Model Configuration Params Activated Params Pile PPL ($\downarrow$)
@@
3.4 Load Balance Loss
LLMs are generally trained in a distributed manner. Yet distributed training can lead to load imbalance of MoE models Lepikhin et al. (2021); Fedus et al. (2022); Dai et al. (2022), where a minority of expert handle the majority of tokens while the others remain idle most of the time. That can negatively impact training efficiency. It is generally desirable for the number of tokens processed by different experts to be roughly equal. To achieve that, a load balancing loss is commonly introduced in the training of MoE models. We follow previous works Huang et al. (2024); Fedus et al. (2022) and adopt a widely used loss as follows:
$\mathcal{L}_{bal}$ $=N\cdot\sum_{i=1}^{N}w_{i}R_{i}$ (7)
[…]
$R_{i}$ $=\frac{1}{B}\sum_{j=1}^{B}\mathbf{r}_{i}^{j}$
where $B$ represents the number of tokens in a mini-batch. $\mathbf{r}^{j}$ denotes the probability distribution of the routing output for the $j$-th token, derived via Eq. 7, and $\mathbf{r}_{i}^{j}$ represents the specific probability value w.r.t the $i$-th expert. It is noteworthy that for infrequent tokens, which have only one visible expert, the load balancing loss does not apply to them. Since infrequent tokens are assigned to experts through uniform random hashing, the overall load is supposed to be natively balanced, thus having no need for the balancing loss. The balancing loss primarily regulates the routing w.r.t frequent tokens.
Our final loss is a combination of the language model loss and the load-balance loss:
$\mathcal{L}=\mathcal{L}_{lm}+\mathcal{L}_{bal}$ (8)
4 Experiments